A given mass of gas occupies a volume of 285 mL at 24°C and 741 mmHg. What will be the new volume at STP?

Volume= mL
Answer #1

At STP! Temperature, T-273 K , Pressure, P= 760 mm Hg - 1 atmIda Gas law! PV = nRT-C) P » Pressure , > Volume N No. of mole, Ra universal gas constant T temperature in Kelvin Modified frGriven, P, = 741 mmHly, V = 285 mL, T, = (24+973)K = 297K P2 = 760 mmtly, z = ?, T2 = 273K. i V = 741 Monty x 285 mL . 2734 2